3questions in total please *Problem 6.66 You stand on a spherical asteroid of uniform density whose mass is 2.5 x 1016 kg and whose radius is 11 km (11Which of the following are true in this problem? The final kinetic energy of the rock is 0 The total energy of the rock+aster
Answer #1

Given :

Mass of asteroid (M) = 2.5 ×1016 Kg

Radius of asteroid (R) = 11× 103 m

Final velocity of rock (V) =10 m/s

Lets mass of rock is 'm'

As there is no external forces are acting So, the energy will be conserve therefore :

T.Ef = T.Ei

Lets final and initial kinetic energy of rock are Kf and Ki respectively

Similarly,

Final and initial potential energy of rock are Pf and Pi respectively.

So,

Ki + Pi = Kf + Pf

GmM

22GAM

2 × 6.674 × 10-11 × 2.5 × 1016 11000 +10

2 × 6.674 × 2.5 × 102 +102 V)

v= 20 m/s

So, initial speed of rock to not com back at asteroid is 20m/s

Now,

(a) As there is some final speed so the final kinetic energy is not ZERO

(b) Total energy of (rock + asteroid will be conserve so in both state Total energy will be EQUAL

(c) initially velocity is 20 m/s and final velocity is 10 m/s so final kinetic energy of rock is LESS than initially kinetic energy of rock

(d) As the velocities are of same order so change in kinetic energy is NOT NEGLIGIBLE

(e) As the final kinetic energy is less So to conserve the Total energy final potential energy will be MORE than the initial potential energy